John and Janet were moving a heavy crate (with a mass M=250 kg) from their leveled driveway into their garage. Janet was pushing the crate horizontally with a force equal to 1/5 her weight. Her mass is 51 kg. John was pulling the crate with a 500N force at a 30 degree angle with the horizontal. The driveway and the wooden crate have a frictional forcve of 50.0N between them.

a.) total work done on the crate when it moved 45 m.
b.) It took 120 seconds to move the crate, calculate power.
1
Expert's answer
2011-06-15T07:12:18-0400
a) Work a=F*s, F=force, s- distance
Total work si& A=A(Janet)+A(John)-A(friction)
Hence we have
A=(51/5*9.81+500*cos(30)-50)*45=21738 Joules

b) Power P=A/t
Hence P=21738/120=181.2 Joules/s=181.2 Watt